Circular No. 7554

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

2000 WR_{106}

D. Jewitt and H. Aussel, University of Hawaii, report their measurement of submillimeter continuum radiation from 2000 WR_{106} (cf. MPEC 2000-X02, 2000-Y45) using the 15-m James Clerk Maxwell Telescope on Mauna Kea: “A preliminary estimate of the flux density at 850 μm on 2000 Dec. 30 and 31 gives 2.5 ± 0.8 mJy. Simultaneous R-band optical photometry by A. Evans (SUNY at Stony Brook) using the University of Hawaii 2.2-m telescope yields red mag 19.7. The submillimeter flux, interpreted as Rayleigh-Jeans thermal emission, corresponds to an object diameter of 900_{−150}^{+100} km. The derived geometric albedo is 7_{−1.5}^{+3} percent. Thus, 2000 WR_{106} is about half the size of Pluto. Its surface is, on average, about eight times darker, suggesting the absence of widespread surface frost.”

SUPERNOVA 2001A IN NGC 4261 M. Modjaz and W. D. Li, University of California at Berkeley; and M.

Schwartz, Cottage Grove, OR, report the discovery of an apparent supernova (mag ∼ 18.9) by LOTOSS (cf. IAUC 7514) on an unfiltered image taken on Jan. 1.5 UT with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 2001A, confirmed on an image taken on 2000 Dec. 29.5 UT (mag ∼ 18.8) and an image taken with the 0.5-m Tenagra III automatic telescope on Dec. 15.5 (mag ∼ 18.4), is located at α = 12^{h}19^{m}23^{s}.01, δ = +5^{o}49′40′′.5 (equinox 2000.0), which is 3′′.1 west and 10′′.7 north of the nucleus of NGC 4261. The photometry and astrometry of the new object were measured after subtracting the combined image of the new observations from the KAIT observations in May 2000 (limiting mag ∼ 21.0), which showed nothing at the position of the supernova.

POSSIBLE NOVA IN PUPPIS S. Nakano, Sumoto, Japan, reports the following precise position for

this variable (cf. IAUC 7552, 7553) by H. Abe (Yatsuka, 0.26-m f/6.0 reflector) from a CCD image taken on Jan. 1.525 UT: α = 7^{h}37^{m}56^{s}.91, δ = −25^{o}56′59′′.7 (equinox 2000.0). G. Masi, Ceccano, Italy, reports the following position end figures from an image taken by M. Uemura (Kyoto University): 56^{s}.91, 59′′.0. Masi notes the presence of a USNO-A2.0 catalogue star (red mag 13.2) at position end figures 56^{s}.88, 59′′.1. T. Kato forwards position end figures obtained by K. Kadota (Ageo, Saitama, 0.18-m f/5.5 reflector) from an image taken on Jan. 1.654: 56^{s}.88, 58′′.9.

2001 January 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green